OpenStudy (anonymous):

5x+4y=-7 y=2x+8 Use substitution to solve the system.

OpenStudy (anonymous):

Since they around solved for y you just plug in 2x+8 for it 5x+4(2x+8)=-7 5x+8x+32=-7 combine like terms and solve

OpenStudy (anonymous):

Y is already given so you just plug it in, or substitute the y, with the equation. So 5x+4y=-7 becomes 5x+4(2x+8)=-7 Then solve the equation.